OpenStudy (anonymous):

If the radius is is 20 inches then what isthe circumference

OpenStudy (anonymous):

How would u do It

OpenStudy (accessdenied):

\[ C = 2\pi r \] Just use r = 20 and and evaluate.

OpenStudy (anonymous):

10?

OpenStudy (accessdenied):

and no i was not making fun of your 'is is' in the topic, i think i actually put 'and and' for no reason. D: C= 2*pi*20 = 40pi

OpenStudy (anonymous):

Someone ese got 3.185

OpenStudy (accessdenied):

well, assuming all information is correct, 40pi (or if you approximate with the pi, its about 125.6, i didnt check a calculator tho)

OpenStudy (anonymous):

C = 2π r = 40 π = 125.7 in ( You're so confused this result with other problem you posted!